# Leveraging Atom Loss Errors in Fault Tolerant Quantum Algorithms

### v1

@@ 1 Introduction
In particular, qubit loss and leakage from computational subspace are dominant noise sources in many hardware systems, and can have a dramatic effect on the practical performance of quantum processors [14, 15]. In the absence of loss correction, all qubits within a QEC code can eventually disappear, destroying the encoded quantum information. Conversely, directly detecting loss provides valuable information about the error location, in contrast with Pauli errors, which are indirectly inferred from syndrome information. Recent work has shown that the rich information provided by so-called erasure detection can in fact improve QEC performance substantially [16, 17, 18, 19, 10, 11, 20, 21, 22, 23, 24, 25]. Similarly, noise bias present in many platforms [14, 4, 9] can also be leveraged to improve performance [26, 27, 28, 29]. These studies motivate developing strategies that incorporate loss detection and harness bias, to improve practical QEC performance.
@@
Before proceeding, we note that loss-to-erasure conversion has been explored at the memory level, using mid-circuit measurement in alkaline-earth-like atomic systems [17, 10, 16, 11] and superconducting qubits [19, 21, 20, 15, 36, 37, 38], as well as using leakage-reduction units with extra qubits and gates [25, 39, 40, 41]. Since the loss must be addressed in all systems and erasure conversion may not be always directly accessible after each gate operation (or, alternatively, may result in additional time and errors), we focus on the cases involving loss-resolving readout as a part of the natural QEC cycles. Most recently, such approaches were explored within measurement-based quantum computing [22], resulting in thresholds that are significantly lower than mid-circuit erasure conversion methods [17, 10, 16]. Meanwhile, atom loss-resolving readout was demonstrated experimentally and its implications for logical qubit performance were explored in Refs. [24, 23, 8]. We extend these previous studies, which focus on the performance of individual logical qubits, to analyze atom loss effects at all levels ranging from decoding and QEC architecture to logical algorithms. We systematically compare different loss detection and correction methods, including their frequencies and associated space-time overheads. By focusing on logical algorithms we provide new insights into how circuit structure influences the algorithmic performance and develop new, optimized strategies for circumventing atom loss.
@@ 2 Detecting and decoding delayed erasures
Figure 2: Delayed-erasure decoder. (a) Illustration of a qubit lifecycle and its usage in the delayed-erasure decoder. From initialization to measurement, each physical qubit can be lost at multiple possible time points, each occurring with a potentially different probability and corresponding syndrome. Upon detection of a qubit loss, the decoder accounts for these possibilities in order to improve the accuracy of the assigned correction. (b) Logical error rate for a logical memory as a function of the number of conventional SE rounds before logical measurement, here with distance $d=5$ and loss errors only with probability $p_{\text{loss}}=1\%$ per entangling gate. The delayed-erasure decoder (pink) substantially outperforms a decoder which does not account for loss information (black). The erasure channel (yellow) gives a lower bound on the possible logical error rate, as its performance is optimal. As lifecycle length increases, the delayed-erasure decoder alone is insufficient to match the erasure channel, and active loss detection and replacement are needed to maintain performance.
@@
Formalizing our approach, our goal is to automatically construct a decoding graph based on the observed loss, which captures how errors (hyperedges) trigger checks (vertices that compare consecutive stabilizer measurements in time). Ideally, we would solve the MLE decoding problem, which identifies the most likely configuration of Pauli and loss errors consistent with both the observed error syndromes and loss events (see Supplementary Materials [49] Section S2.1). However, since a full MLE solution would require considering all combinations of loss locations due to the non-additive nature of loss-induced errors, this approach is computationally intractable for even modest system sizes and low loss rates.
@@
Instead, we approximate the MLE solution by handling each loss event independently (see Supplementary Materials [49] Section S2.2 for full details). For each loss event, we trace back the qubit’s lifecycle, accounting for all potential loss events and their associated probabilities. Each potential loss event corresponds to a loss circuit, in which certain gates are canceled due to the loss. Using Stim [52], a Clifford circuit simulator, we construct the hyperedges and their probabilities for each loss circuit. Next, we integrate all hyperedges within a qubit’s lifecycle, re-weighting them based on the probabilities of their corresponding potential loss locations. In cases of multiple losses, each loss event is calculated independently, and the results are combined to construct the final decoding graph, which serves as the input to the decoder. Empirically, this heuristic achieves performance comparable to methods that consider combinations of loss events, while significantly reducing computational overhead.
@@
To benchmark the performance of our decoder, we focus on the well-studied case of a logical memory experiment on a surface code, which involves repeated SE rounds over time, while loss is detected only by SSR. In this setting, losses on measure qubits are detected every SE round, but losses on data qubits are not detected until the logical qubit is projectively measured at the end of the circuit. We perform simulations on a single surface code with code distance $d=5$ and circuit-level noise with an entangling gate loss error of $1\%$. Unless otherwise noted, this paper considers loss and Pauli errors occurring during entangling operations, as this represents a dominant source of error in many quantum computing architectures [4, 9, 14]. The error model details are in the Supplementary Materials Section S5.2.3, with a comparison to other error models  [49].
@@
Figure  2(b) presents the improvement in logical error rate for the delayed-erasure decoder (pink lines) over an MLE decoder (black lines) which does not account for loss information. The delayed-erasure decoder improves the logical error rate by several orders of magnitude. Furthermore, for shallower circuits ($\lesssim 2$ SE rounds), the delayed-erasure performance is close to the optimal performance corresponding to an erasure channel (yellow lines), where loss is detected and corrected immediately, corresponding to a lifecycle length of one. This is because in shallower circuits, the propagation of loss errors is nearly deterministic, greatly simplifying the decoding problem and enabling the delayed-erasure decoder to effectively handle these errors.
@@
In contrast, when the circuit depth is large prior to logical measurement, the qubit lifecycles are extended and loss errors can occur at multiple points. These losses introduce correlated errors, depending on when and where they occur. While the delayed-erasure decoder performs relatively well in this setting compared to a decoder which does not utilize loss information (as shown in Fig. 2(b)), its effectiveness deteriorates compared to the erasure channel when lifecycles become too long due to error accumulation. To address this, loss detection and qubit replacement must be incorporated into the algorithm to maintain performance. Therefore, in the next section we explore practical approaches to detect and replace lost atoms with minimal overhead.
@@ 3.1 Modified conventional SE
Each SWAP operation incurs a cost due to qubit movement, making it crucial to evaluate different SWAP periods to balance performance and experimental complexity. The SWAP period defines how often loss-detecting SE rounds are interspersed with conventional SE rounds, occurring at a fixed interval. We identify a trade-off between the lifecycle lengths of data and measurement qubits for different SWAP periods, demonstrating that a period of 2 can be competitive with a period of 1 (see Supplementary Materials [49] Figs. S3, S4). A period of 2 reduces experimental complexity while maintaining a similar average lifecycle length.
@@ 3.2 Teleportation-based SE
Teleportation-based SE is an alternative QEC approach which utilizes regular teleportation between logical qubit layers, inherently enabling loss detection upon logical measurement (see Fig. 3(b)). While typically associated with measurement-based quantum computation (MBQC) architectures, the teleportation-based SE can be used in a circuit-based framework where gates are directly applied on qubits. We realize the XZZX cluster state [53, 54, 29], as described in detail in the Supplementary Materials [49] Section S3.5. In this framework, each qubit is entangled with four neighboring qubits, which, under our error model considering noise on entangling gates only, results in short lifecycles of length four. Among the SSR-based approaches considered, this SE method enables the shortest lifecycles with no experimental overhead apart from SSR, though it requires additional qubits. These features make it well-suited for neutral atom architectures and particularly effective in loss-dominant regimes, though its advantages diminish when the loss fraction is small, as detailed in Section 3.4.
@@ 3.3 Modified Steane SE
Figure 3: Loss-detecting SE methods. (a) Modified SWAP SE method for detecting and correcting loss errors using physical SWAP operations between data and measure qubits, without gate overhead. (b) Teleportation-based SE method, utilizing layers of teleportation and stabilizer checks in alternating bases. The circuit is repeated $d$ times. (c) Steane SE method with logical SWAP, employing logical fault-tolerant measurement qubits and teleportation. (d) Logical error rates as a function of SE rounds using the delayed-erasure decoder, with a code distance of 7 and a physical error rate $1\%$. Comparable performance is observed across SE methods in regimes with short lifecycles. The erasure channel, realized through Free SE with period 0.25, provides a lower bound. (e) Error thresholds as a function of the loss fraction for different SE methods, showcasing the improvement with loss for all methods. (f) Effective distance as a function of loss fraction for different SE methods, considering distance $d=7$. (g) Space-time overhead as a function of the loss fraction, describing the required distance to achieve logical error rates of $10^{-12}$ given a physical error rate of $0.5\%$. The legend for (d-g) appears in (e).
@@ 3.4 Comparison of SE methods
We also evaluate a Free SE method, which detects and replaces lost qubits at different intervals: after every gate (period 0.25) or every $K$ rounds (period $K$). A possible implementation of this approach involves direct erasure conversion and atom replacement [17]. Free detection at period 0.25 eliminates loss propagation, ensuring that the exact loss locations are known, effectively realizing an erasure channel. This sets a lower bound for the possible logical error rates.
Figure 3(d) compares the logical error rates of conventional SE without loss detection, approaches with loss detection—such as SWAP SE, teleportation-based SE, and Free SE—and the erasure channel in a single logical memory experiment at an experimentally-motivated loss fraction of 0.5. For a small number of SE rounds ($\leq 10$), all protocols behave similarly. However, as the number of SE rounds increases, loss detection and the replacement of lost qubits substantially enhance performance, providing a convenient approach to achieving high circuit depths without additional experimental gate overhead.
@@
Figure 3(f) presents the effective code distance of each SE method as a function of the loss fraction, for distance $d=7$. The effective distance $d_{e}$ refers to the number of errors required to cause a logical failure, as the logical error rate scales as $(p/p_{\text{th}})^{d_{e}}$ far below the threshold $p_{\text{th}}$, where $p$ is the physical error rate. For Pauli noise, $d_{e}=(d+1)/2$, and for erasure noise $d_{e}=d$. To determine $d_{e}$ for each loss fraction, we fit the logical error rate data far below the threshold to the function $\alpha p^{\beta}$, where $\alpha$ and $\beta$ are fitting parameters (see Supplementary Materials [49] Fig. S13). As expected, all SE methods, while utilizing the delayed-erasure decoder, experience increased effective distance with an increased loss fraction. For loss errors only, teleportation-based SE and Free SE methods (periods 0.25 and 1) achieve the optimal effective distance of $d_{e}\approx d$, while SWAP SE achieves $d_{e}\approx d-1$, likely due to longer lifecycles.
@@
Another important aspect of neutral atom qubits is their intrinsic bias, as $Z$-type Pauli errors are much more common than $X$-type Pauli errors [14]. In Supplementary Materials Section S5.2.2 [49], we investigate the interplay between biased errors and delayed erasure errors across multiple SE methods, employing the XZZX surface code [27] and the XZZX cluster state [29] for teleportation-based SE. Our analysis considers scenarios both with and without bias-preserving gates, as well as the presence of biased erasure noise. Bias-preserving gates, such as native $CX$ gates, preserve the bias of the errors, in contrast to $CX$ gates decomposed into $CZ$ and $H$ gates. Biased erasure noise occurs when a qubit exits the computational subspace exclusively from the state |$\ket{1}$. As a result, the replaced qubit follows a biased error channel instead of a depolarizing error channel. Additionally, we investigate scenarios where Pauli errors are biased, examining the two-dimensional space defined by the bias ratio and the loss fraction. Thresholds are calculated at each point in this space for various SE methods (see Figures S14 and S15 in the Supplementary Materials [49]). A key observation emerges across all SE methods: increasing the loss fraction has a significantly greater impact on thresholds than increasing the bias ratio, even when using bias-preserving gates and biased-erasure. This difference in impact between the loss fraction and the bias ratio becomes even more pronounced in the absence of bias-preserving gates.
@@ 3.5 Predicting performance by error counting
The curve intersections with the axes provide key insights. The y-axis intersection, representing the threshold in the absence of Pauli errors, is related to the lifecycle length. For the erasure channel, achieved using the Free SE with a period of 0.25, the threshold is $p_{\text{loss,th}}=6.9\%$. This is consistent with the 3D bond percolation threshold of $25\%$, scaled by the four gates per qubit used in circuit-level simulations [47, 63, 64]. Increasing the lifecycle reduces the threshold, which decays with the lifecycle length, as shown in Fig. 4(b) (see Supplementary Materials [49] Section S3.7 for further details). Notably, the SWAP SE, with an average lifecycle length of 8, has a lower threshold of $p_{\text{loss,th}}=2\%$, deviating from the heuristic. This lower performance is attributed to additional factors unique to SWAP SE, such as the inability to detect measurement qubit loss when data qubits are detected as lost and the loss of stabilizer information, as discussed previously in Section 3.1.

### v2

@@ 1 Introduction
In particular, qubit loss and leakage from computational subspace are dominant noise sources in many hardware systems, and can have a dramatic effect on the practical performance of quantum processors [14, 15]. In the absence of loss correction, all qubits within a QEC code can eventually disappear, destroying the encoded quantum information. Conversely, directly detecting loss provides valuable information about the error location, in contrast with Pauli errors, which are indirectly inferred from syndrome information. Recent work has shown that the rich information provided by so-called erasure detection can in fact improve QEC performance substantially [16, 17, 18, 19, 10, 11, 20, 21, 22, 23, 24, 25, 26]. Similarly, noise bias present in many platforms [14, 4, 9] can also be leveraged to improve performance [27, 28, 29, 30]. These studies motivate developing strategies that incorporate loss detection and harness bias, to improve practical QEC performance.
@@
Before proceeding, we note that loss-to-erasure conversion has been explored at the memory level, using mid-circuit measurement in alkaline-earth-like atomic systems [17, 10, 16, 11] and superconducting qubits [19, 21, 20, 15, 37, 38, 39], as well as using leakage-reduction units with extra qubits and gates [25, 40, 41, 42]. Since the loss must be addressed in all systems and erasure conversion may not be always directly accessible after each gate operation (or, alternatively, may result in additional time and errors), we focus on the cases involving loss-resolving readout as a part of the natural QEC cycles. Most recently, atom loss-resolving readout was demonstrated experimentally and its implications for logical qubit performance were explored in Refs. [22, 24, 23, 8]. We extend these previous studies, which focus on the performance of individual logical qubits, to analyze atom loss effects at all levels ranging from decoding and QEC architecture to logical algorithms. We systematically compare different loss detection and correction methods, including their frequencies and associated space-time overheads. By focusing on logical algorithms we provide new insights into how circuit structure influences the algorithmic performance and develop new, optimized strategies for circumventing atom loss.
@@ 2 Detecting and decoding delayed erasures
Figure 2: Delayed-erasure decoder. (a) Illustration of a qubit lifecycle and its usage in the delayed-erasure decoder. From initialization to measurement, each physical qubit can be lost at multiple possible time points, each occurring with a potentially different probability and corresponding syndrome. Upon detection of a qubit loss, the decoder accounts for these possibilities in order to improve the accuracy of the assigned correction. (b) Logical error rate for a logical memory as a function of the number of conventional SE rounds before logical measurement, here with distance $d=5$ and loss errors only with probability $p_{\text{loss}}=1\%$ per entangling gate. The delayed-erasure decoder (pink) substantially outperforms a decoder which does not account for loss information (black). Even without accurate loss information, the delayed-erasure decoder achieves comparable performance to a decoder with perfect loss information (gray, dashed).
@@
Formalizing our approach, our goal is to automatically construct a decoding graph based on the observed loss, which captures how errors (hyperedges) trigger checks (vertices that compare consecutive stabilizer measurements in time). Ideally, we would solve the MLE decoding problem, which identifies the most likely configuration of Pauli and loss errors consistent with both the observed error syndromes and loss events (see Appendix Section BB.1). However, since a full MLE solution would require considering all combinations of loss locations due to the non-additive nature of loss-induced errors, this approach is computationally intractable for even modest system sizes and low loss rates.
@@
Instead, we approximate the MLE solution by handling each loss event independently (see Appendix Section BB.2 for full details). For each loss event, we trace back the qubit’s lifecycle, accounting for all potential loss events and their associated probabilities. Each potential loss event corresponds to a loss circuit, in which certain gates are canceled due to the loss. Using Stim [53], a Clifford circuit simulator, we construct the hyperedges and their probabilities for each loss circuit. Next, we integrate all hyperedges within a qubit’s lifecycle, re-weighting them based on the probabilities of their corresponding potential loss locations. In cases of multiple losses, each loss event is calculated independently, and the results are combined to construct the final decoding graph, which serves as the input to the decoder. Empirically, this heuristic achieves performance comparable to methods that consider combinations of loss events, while significantly reducing computational overhead.
@@
To benchmark the performance of our decoder, we focus on the well-studied case of a logical memory experiment on a surface code, which involves repeated SE rounds over time, while loss is detected only by SSR. In this setting, losses on measure qubits are detected every SE round, but losses on data qubits are not detected until the logical qubit is projectively measured at the end of the circuit. We perform simulations on a single surface code with code distance $d=5$ and circuit-level noise with an entangling gate loss error of $1\%$. Unless otherwise noted, this paper considers loss and Pauli errors occurring during entangling operations, as this represents a dominant source of error in many quantum computing architectures [4, 9, 14]. Details of the error model, along with results using alternative error models, can be found in Supplementary Materials Section S4.2.3 [50] and Table 1.
@@
We begin by analyzing the case without any loss replacement, while increasing the number of rounds—and therefore, the lifecycle length. Figure 2(b) shows the logical error rate for three types of decoders: the MLE decoder, which does not account for loss information (black lines); the delayed-erasure decoder, which uses SSR information (pink lines); and an erasure decoder with perfect loss location information (gray dashed lines). Despite lacking accurate loss information, the delayed-erasure decoder improves the logical error rate by several orders of magnitude. Furthermore, even for deep circuits with long lifecycles, its performance is comparable to that of decoding with perfect loss information. As we will see later, for shallower circuits, the performance of the delayed-erasure decoder approaches that achieved with both perfect loss location information and full lost qubit replacement every round. This is because, in shallow circuits, the propagation of loss errors is nearly deterministic, significantly simplifying the decoding problem and allowing the delayed-erasure decoder to effectively handle these errors.
@@
In contrast, when the circuit depth is large prior to logical measurement, the qubit lifecycles are extended and loss errors can occur at multiple points. These losses introduce correlated errors, depending on when and where they occur. While the delayed-erasure decoder performs relatively well in this setting compared to a decoder which does not utilize loss information and also compared to the decoder with perfect loss information, its effectiveness deteriorates when lifecycles become too long due to error accumulation. To address this, loss detection and qubit replacement must be incorporated into the algorithm to maintain performance. Therefore, in the next section we explore practical approaches to detect and replace lost atoms with minimal overhead.
@@ 3.1 Modified conventional SE
Each SWAP operation incurs a cost due to qubit movement, making it crucial to evaluate different SWAP periods to balance performance and experimental complexity. The SWAP period defines how often loss-detecting SE rounds are interspersed with conventional SE rounds, occurring at a fixed interval. We identify a trade-off between the lifecycle lengths of data and measurement qubits for different SWAP periods, demonstrating that a period of 2 can be competitive with a period of 1 (see Supplementary Materials [50] Figs. S2, S3). A period of 2 reduces experimental complexity while maintaining a similar average lifecycle length.
@@ 3.2 Teleportation-based SE
Teleportation-based SE is an alternative QEC approach which utilizes regular teleportation between logical qubit layers, inherently enabling loss detection upon logical measurement [54, 55, 22, 16] (see Fig. 3(b)). While typically associated with measurement-based quantum computation (MBQC) architectures, the teleportation-based SE can be used in a circuit-based framework where gates are directly applied on qubits. We realize the XZZX cluster state [30], as described in detail in the Supplementary Materials [50] Section S2.5. In this framework, each qubit is entangled with four neighboring qubits, which, under our error model considering noise on entangling gates only, results in short lifecycles of length four. Among the SSR-based approaches considered, this SE method enables the shortest lifecycles with no experimental overhead apart from SSR, though it requires additional qubits. These features make it well-suited for neutral atom architectures and particularly effective in loss-dominant regimes, though its advantages diminish when the loss fraction is small, as detailed in Section 3.4.
@@ 3.3 Modified Steane SE
Figure 3: Loss-detecting SE methods. (a) Modified SWAP SE method for detecting and correcting loss errors using physical SWAP operations between data and measure qubits, without gate overhead. (b) Teleportation-based SE method, utilizing layers of teleportation and stabilizer checks in alternating bases. The circuit is repeated $d$ times. (c) Steane SE method with logical SWAP, employing logical fault-tolerant measurement qubits and teleportation. (d) Logical error rates as a function of SE rounds using the delayed-erasure decoder, with a code distance of 7 and a physical error rate $1\%$. Comparable performance is observed across SE methods in regimes with short lifecycles. (e) Error thresholds as a function of the loss fraction for different SE methods, showcasing the improvement with loss for all methods. (f) Effective distance as a function of loss fraction for different SE methods, considering distance $d=7$. (g) Space-time overhead as a function of the loss fraction, describing the required distance to achieve logical error rates of $10^{-12}$ given a physical error rate of $0.5\%$. The legend for (d-g) appears in (e).
@@ 3.4 Comparison of SE methods
We also evaluate a Free SE method, which detects and replaces lost qubits at different intervals: after every gate (period 0.25) or every $K$ rounds (period $K$). Free detection at period 0.25 eliminates loss propagation, ensuring that the exact loss locations are known, effectively realizing an erasure channel. Furthermore, we compare our results to a Free SE method with period 1, including an additional feature: the exact loss locations are known, greatly simplifying the decoding problem. A possible implementation of these SE approaches involves mid-circuit erasure conversion and atom replacement [17].
Figure 3(d) compares the logical error rates of conventional SE without loss detection, approaches with loss detection—such as SWAP SE, teleportation-based SE, and Free SE—and Free SE with perfect loss information, in a single logical memory experiment at an experimentally-motivated loss fraction of 0.5. For a small number of SE rounds ($\leq 10$), all protocols behave similarly. However, as the number of SE rounds increases, loss detection and the replacement of lost qubits substantially enhance performance, providing a convenient approach to achieving high circuit depths without additional experimental gate overhead.
@@
Figure 3(f) presents the effective code distance of each SE method as a function of the loss fraction, for distance $d=7$. The effective distance $d_{e}$ refers to the number of errors required to cause a logical failure, as the logical error rate scales as $(p/p_{\text{th}})^{d_{e}}$ far below the threshold $p_{\text{th}}$, where $p$ is the physical error rate. For Pauli noise, $d_{e}=(d+1)/2$, and for erasure noise $d_{e}=d$. To determine $d_{e}$ for each loss fraction, we fit the logical error rate data far below the threshold to the function $\alpha p^{\beta}$, where $\alpha$ and $\beta$ are fitting parameters (see Supplementary Materials [50] Fig. S13). As expected, all SE methods, while utilizing the delayed-erasure decoder, experience increased effective distance with an increased loss fraction. For loss errors only, teleportation-based SE and Free SE methods (periods 0.25 and 1) achieve the optimal effective distance of $d_{e}\approx d$, while SWAP SE achieves $d_{e}\approx d-1$, likely due to longer lifecycles. Additional threshold and effective distance results under an alternative error model from Ref. [22] are presented in Table 1.
@@
Another important aspect of neutral atom qubits is their intrinsic bias, as $Z$-type Pauli errors are much more common than $X$-type Pauli errors [14]. In Supplementary Materials Section S4.2.2 [50], we investigate the interplay between biased errors and delayed erasure errors across multiple SE methods, employing the XZZX surface code [28] and the XZZX cluster state [30] for teleportation-based SE. Our analysis considers scenarios both with and without bias-preserving gates, as well as the presence of biased erasure noise. Bias-preserving gates, such as native $CX$ gates, preserve the bias of the errors, in contrast to $CX$ gates decomposed into $CZ$ and $H$ gates. Biased erasure noise occurs when a qubit exits the computational subspace exclusively from the state |$\ket{1}$. As a result, the replaced qubit follows a biased error channel instead of a depolarizing error channel. Additionally, we investigate scenarios where Pauli errors are biased, examining the two-dimensional space defined by the bias ratio and the loss fraction. Thresholds are calculated at each point in this space for various SE methods (see Figures S16 and S17 in the Supplementary Materials [50]). A key observation emerges across all SE methods: increasing the loss fraction has a significantly greater impact on thresholds than increasing the bias ratio, even when using bias-preserving gates and biased-erasure. This difference in impact between the loss fraction and the bias ratio becomes even more pronounced in the absence of bias-preserving gates.
@@ 3.5 Predicting performance by error counting
The curve intersections with the axes provide key insights. The y-axis intersection, representing the threshold in the absence of Pauli errors, is related to the lifecycle length. For the Free SE period 1 with perfect loss information, the threshold is $p_{\text{loss,th}}\approx 7.2\%$. Increasing the lifecycle reduces the threshold, which decays with the lifecycle length, as shown in Fig. 4(b) (see Supplementary Materials [50] Section S2.7 for further details). Notably, the SWAP SE, with an average lifecycle length of 8, has a lower threshold of $p_{\text{loss,th}}\approx 2.3\%$, deviating from the heuristic. This lower performance is attributed to additional factors unique to SWAP SE, such as the inability to detect measurement qubit loss when data qubits are detected as lost and the loss of stabilizer information, as discussed previously in Section 3.1.